OpenStudy (anonymous):

Please help:) write the first five terms of the geometric sequences.. I suck at this okay so its... A1= 1, R= 1/3

OpenStudy (anonymous):

apply formula an = a1r ^(n -1) for each

OpenStudy (anonymous):

\(a_1=1,a_2=1\times \frac{1}{3}=\frac{1}{3},a_3=\frac{1}{3}\times \frac{1}{3}=\frac{1}{9}\) etc

OpenStudy (anonymous):

thanks